Direct material cost  = £90 

Direct labour cost = £135 

Other variable costs = £185 

Fixed overhead cost = 25,000 

Care home capacity = 45 clients 

SALES per year = £11,500 per annum


a) Determine the break-even capacity usage rate for ABC Care Home ltd. 

(b) Calculate the targeted profit if the care home achieves 90% capacity usage.


1
Expert's answer
2022-03-14T18:05:05-0400

Break-Even point (units) = Fixed Costs ÷ (Sales price per unit – Variable costs per unit)

Variable cost per unit= "\frac{90+135+185}{45}" = 9

Sales price per unit="\frac{11500}{45}= 256"

BEP="\frac{25000}{256-9}= 101" clients

b) Targeted profit

90% capacity= 0.9×145= 130.5 clients

TCM=230.5× 247= 32,233.5

Profit= 32,233.5-25000= 7,233.5